Solve: 1/2x = 17
I need the answer to this.

Answer by nerdybill(5404)

There is some uncertainty in the way you wrote your equation.
There are two ways I could read it:
1/(2x) = 17
or
(1/2)x = 17
.
I'll assume the latter. To solve, simply isolate the unknown (x) to one side of the equation -- doesn't matter which side.
(1/2)x = 17
multiply both sides by 2:
x = 34 (this is your solution)